# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v3

@@ V-A Bayesian Inference
Since the measurement matrix $\boldsymbol{\Psi}$ may be “bad” (e.g., rank-deficient, ill-conditioned, or having a non-zero mean) [39], the divergence issues might be arisen in the Bayesian inference. To address this, we first perform unitary transformations on the measurement matrix, i.e., $\boldsymbol{\Psi}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$. Consequently, the original signal model in (30) is rewritten as
@@
$\hat{x}_{n,k}=\int{x}_{n,k}p(x_{n,k}|\mathbf{R})\mathrm{d}x_{n,k},$ (37)
where the posterior distribution $p(x_{n,k}|\mathbf{R})$ is given by
$\int p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\,\mathrm{d}\beta\,\mathrm{d}\mathbf{Z}\,\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\,\mathrm{d}\mathbf{\boldsymbol{\Omega}}\,\mathrm{d}\mathbf{S}\,\mathrm{d}\mathbf{X}_{\backslash n,k}$ (38)
with $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ indicating the joint posterior probability and satisfying
$p(\beta,\mathbf{X},\mathbf{Z},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\propto p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S},\mathbf{R})$ (39)
$=$ $p(\mathbf{R}|\mathbf{Z},\beta)p(\mathbf{Z}|\mathbf{X})p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S})p(\beta),$
where $p(\beta)\propto\beta^{-1}$ denotes the noise precise level; $p(\mathbf{R}|\mathbf{Z},\beta)$ and $p(\mathbf{Z}|\mathbf{X})$ are given by
$p(\mathbf{R}|\mathbf{Z},\beta)$ $=\frac{1}{(\pi\beta^{-1})^{MK}}\prod_{m}\prod_{k}p(r_{m,k}\mid z_{m,k}),$ (40)
@@ V-B Factor Graph Representation
The dependencies among the random variables in the factorization (39) are illustrated using a factor graph, as shown in Fig. 3, with the corresponding probability distributions summarized in Table I. It is evident that the factor graph in Fig. 3 is significantly more intricate than those considered in [39, 43, 40, 42, 41], primarily due to the incorporation of precision parameter sharing mechanism and MRF structure. Specifically, the works in [39, 43, 40] adopt independent priors on the variances of coefficients, thereby promoting element-wise sparsity. However, such models fail to capture the angular or delay-domain correlations among the coefficients. Furthermore, the model in [42] does not include a precision parameter sharing mechanism, and thus overlooks higher-order dependencies among variable nodes. Due to these structural distinctions, existing message passing algorithms developed in [39, 43, 40, 42, 41] are not directly applicable to our setting. Consequently, the message update equations at the variable nodes must be carefully reformulated to accommodate the enhanced complexity of our factor graph.
Figure 3: Factor graph representation for the factorization (39).
@@
Factor Distribution Function
$f_{\beta}$ $p(\beta)$ $\beta^{-1}$
$f_{r_{m,k}}$ $p(r_{m,k}|z_{m,k},\beta)$ $\mathcal{CN}(r_{m,k};z_{m,k},\beta^{-1})$
$f_{z_{m,k}}$ $p(z_{m,k}|\mathbf{x}_{k})$ $\delta(z_{m,k}-\mathbf{A}_{m,:}\mathbf{x}_{k})$
@@
Moreover, in the proposed sparse prior model, the latent precision parameters $\beta$ and $\alpha^{i}_{k}$ follow non-Gaussian distributions. Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure belief propagation (BP) framework.
To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and sum-produce algorithm (SPA) rules. Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying sum-product or variational message passing rules to different edges to enable more flexible and efficient message updates.
The MMV-HMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (37). In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig. 3.
@@ V-C Measurement Module
$m_{f_{r_{m,k}}\rightarrow\beta}$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (42)
@@
$m_{f_{r_{m,k}}\rightarrow z_{m,k}}\propto$ $\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (45)
@@
where $(a)$ is obtained by utilizing (44). Owing to the Gaussian form of the message in (45), we can obtain the following model $\mathbf{r}_{k}=\mathbf{z}_{k}+\boldsymbol{\xi}_{k}$, where $\boldsymbol{\xi}_{k}\sim\mathcal{CN}(\boldsymbol{\xi}_{k};0,\hat{\beta}^{-1}\mathbf{I}_{M\times M})$. This representation facilitates seamless integration with the forward passing of UAMP in the measurement module.
Denote the posterior mean and variance of $\mathbf{x}_{k}$ as $\hat{\mathbf{x}}_{k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, which is defined in (64) and (65). According to the UAMP rules [40], define $\boldsymbol{\lambda}$ as $\left|\mathbf{A}\right|^{2}\mathbf{1}_{N}$. Thus, the message from $\mathbf{z}_{k}$ to $f_{\mathbf{r}_{k}}$ is given by $\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})$ with
@@
In this case, utilizing SPA rule, the belief of $\mathbf{z}_{k}$ is given by
$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}$ (48)
$=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M\times M})$
$\overset{(a)}{\propto}\mathcal{CN}(\mathbf{z}_{k};\hat{\mathbf{z}}_{k},\boldsymbol{\nu}_{\mathbf{z}_{k}}),$
where $(a)$ is obtained by the property of Gaussian distribution, i.e., $\mathcal{CN}(x;\mu_{1},\nu_{1})\mathcal{CN}(x;\mu_{2},\nu_{2})\propto\mathcal{CN}(x;\mu,\nu)$ with $\nu=(1/\nu_{1}+1/\nu_{2})^{-1}$ and $\mu=\nu(\mu_{1}/\nu_{1}+\mu_{2}/\nu_{2})$; In addition, $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ are respectively given by
@@ V-D1 Forward Message Passing
$m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (52)
@@
Furthermore, according to SPA rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
@@
where $(a)$ is obtained by utilizing the approximation $b(\boldsymbol{\alpha}_{k})=b(\alpha^{1}_{k})b(\alpha^{2}_{k})\approx\prod_{j\neq n}m_{\boldsymbol{\alpha}_{k}\rightarrow f_{\gamma_{j,k}}}$ and the expectation of Gamma distribution; $\pi^{\mathrm{out}}_{n}$ is defined as
@@
Utilizing the message $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})$, we further derive the message update in the MRF with a 4-connect scheme, where $s_{n,k}^{{\text{l}}}\triangleq s_{n,k-1}$, $s_{n,k}^{\text{r}}\triangleq s_{n,k+1}$, $s_{n,k}^{{\text{t}}}\triangleq s_{n-1,k}$, and $s_{n,k}^{{\text{b}}}\triangleq s_{n,k+1}$ denote the left, right, top and bottom neighbors of $x_{n,k}$. The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions. Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SPA rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (54), as shown in the top of this page.
@@ V-D2 Backward Message Passing
With the messages of neighbors and $v(s_{n,k})$ and SPA rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
@@
According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
@@
where $(a)$ is obtained by utilizing the property of delta function, i.e., $\int\delta(x-t)f(x)\mathrm{d}x=f(t)$. Similarly, $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{2}}\propto\alpha_{k}^{2}\exp(-\alpha_{k}^{2}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$. As a result, the belief of $\alpha_{k}^{1}$ is given by
${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (58)
where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (59)
@@
$b(\gamma_{n,k})\propto m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})$ (60)
@@
$m_{f_{x_{n,k}}\rightarrow x_{n,k}}$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (61)

### v4

@@ V-A Bayesian Inference
Since the measurement matrix $\boldsymbol{\Psi}$ may be “bad” (e.g., rank-deficient, ill-conditioned, or having a non-zero mean) [39], the divergence issues might be arisen in the Bayesian inference. To address this, we first perform unitary transformations on the received signal. Let the singular value decomposition (SVD) of the measurement matrix be denoted as $\boldsymbol{\Psi}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{V}^{\mathrm{H}}$, where $\mathbf{U}$ and $\mathbf{V}$ are two unitary matrices. Performing a unitary transformation with $\mathbf{U}^{\mathrm{H}}$ on the received signal model in (30) yields the following model
@@
$\hat{x}_{n,k}=\int{x}_{n,k}p(\boldsymbol{\Theta}|\mathbf{R})\mathrm{d}\beta\mathrm{d}\mathbf{Z}\mathrm{d}\mathbf{X}\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\mathrm{d}\mathbf{\boldsymbol{\Omega}}\mathrm{d}\mathbf{S},$ (37)
where $\boldsymbol{\Theta}\triangleq\{\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}\}$, and $p(\boldsymbol{\Theta}|\mathbf{R})$ denotes the joint posterior probability, which is defined as
$p(\boldsymbol{\Theta}|\mathbf{R})\propto p(\mathbf{R}|\mathbf{Z},\beta)p(\mathbf{Z}|\mathbf{X})p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})p(\beta),$ (38)
where $p(\beta)\propto\beta^{-1}$ denotes the prior distribution of noise precision; the conditional distributions $p(\mathbf{R}|\mathbf{Z},\beta)$ and $p(\mathbf{Z}|\mathbf{X})$ are respectively given by
$p(\mathbf{R}|\mathbf{Z},\beta)$ $=\prod_{m}\prod_{k}p(r_{m,k}\mid z_{m,k}),$ (39)
@@ V-B Factor Graph Representation
Factor
Distribution
Function
$f_{\beta}$ $p(\beta)$
$\beta^{-1}$
$f_{r_{m,k}}$ $p(r_{m,k}|z_{m,k},\beta)$
$\mathcal{CN}(r_{m,k};z_{m,k},\beta^{-1})$
$f_{z_{m,k}}$ $p(z_{m,k}|\mathbf{x}_{k})$
$\delta(z_{m,k}-\mathbf{A}_{m,:}\mathbf{x}_{k})$
@@
The dependencies among the random variables in the factorization (38) are illustrated in Fig. 3, with the corresponding probability distributions summarized in Table I. It is evident that the factor graph in Fig. 3 is significantly more intricate than those considered in [39, 43, 40, 42, 41], primarily due to the incorporation of precision parameter sharing mechanism and MRF structure. Specifically, the works in [39, 43, 40] adopt independent priors on the variances of coefficients, thereby promoting element-wise sparsity. However, such models fail to capture the angular or delay-domain correlations among the coefficients. Furthermore, the model in [42] does not include a precision parameter sharing mechanism, and thus overlooks higher-order dependencies among variable nodes. Due to these structural distinctions, existing message passing algorithms developed in [39, 43, 40, 42, 41] are not directly applicable to our setting. Consequently, the message update equations at the variable nodes must be carefully reformulated to accommodate the enhanced complexity of our factor graph.
Figure 3: Factor graph representation for the factorization (38).
Moreover, in the proposed sparse prior model, the latent precision parameters $\beta$ and $\alpha^{i}_{k}$ follow non-Gaussian distributions. Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure SP framework.
To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and SP rules. Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying SP or VMP rules to different edges to enable more flexible and efficient message updates.
The MMV-HMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (37). In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules.
@@ V-C Measurement Module
$m_{f_{r_{m,k}}\rightarrow\beta}(\beta)$ $\propto\exp\left\{\int\ln f_{r_{m,k}}b(z_{m,k})\mathrm{d}z_{m,k}\right\},$ (41)
@@
$m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})\propto\exp\left\{\int b(\beta)\ln f_{r_{m,k}}\mathrm{d}\beta\right\}$ (44)
@@
where $(a)$ is obtained by utilizing (43). Owing to the Gaussian form of the message in (44), we can obtain the following model $\mathbf{r}_{k}=\mathbf{z}_{k}+\boldsymbol{\xi}_{k}$, where $\boldsymbol{\xi}_{k}\sim\mathcal{CN}(\boldsymbol{\xi}_{k};0,\hat{\beta}^{-1}\mathbf{I}_{M})$, $\mathbf{r}_{k}=[r_{1,k},r_{2,k},\cdots,r_{M,k}]^{\mathrm{T}}$, and $\mathbf{z}_{k}=[z_{1,k},z_{2,k},\cdots,z_{M,k}]^{\mathrm{T}}$. This representation enables seamless integration with the measurement module of unitary AMP (UAMP) algorithm in [40].
Denote the posterior mean and variance of $\mathbf{x}_{k}$ as $\hat{\mathbf{x}}_{k}$ and $\hat{\nu}_{\mathbf{x}_{k}}$, which is defined in (63) and (64). According to the UAMP algorithm [40], define $\boldsymbol{\lambda}$ as $\left|\mathbf{A}\right|^{2}\mathbf{1}_{N}$. Thus, the message from $\mathbf{z}_{k}$ to $f_{\mathbf{r}_{k}}$ is given by $\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})$ with
@@
In this case, utilizing SP rule, the belief of $\mathbf{z}_{k}$ is given by
$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})$ (47)
$=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\mathcal{CN}(\mathbf{z}_{k};\mathbf{r}_{k},\hat{\beta}^{-1}\mathbf{I}_{M})$
${\propto}\mathcal{CN}(\mathbf{z}_{k};\hat{\mathbf{z}}_{k},\boldsymbol{\nu}_{\mathbf{z}_{k}}),$
where $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\nu}_{\boldsymbol{\mu}_{k}}$ are respectively given by
@@ V-D1 Forward Message Passing
$m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})\propto$ $\exp\left\{\int{b}(x_{n,k})\ln f_{x_{n,k}}\mathrm{d}x_{n,k}\right\}$ (51)
@@
Furthermore, according to SP rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
@@
where $(a)$ is obtained by utilizing the approximation $b(\boldsymbol{\alpha}_{k})=b(\alpha^{1}_{k})b(\alpha^{2}_{k})\approx\prod_{j\neq n}m_{\boldsymbol{\alpha}_{k}\rightarrow f_{\gamma_{j,k}}}(\gamma_{j,k})$ and the expectation of Gamma distribution; $\pi^{\mathrm{out}}_{n}$ is defined as
@@
Utilizing the message $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})$, we further derive the message update in the MRF with a 4-connect scheme, where $s_{n,k}^{{\text{l}}}\triangleq s_{n,k-1}$, $s_{n,k}^{\text{r}}\triangleq s_{n,k+1}$, $s_{n,k}^{{\text{t}}}\triangleq s_{n-1,k}$, and $s_{n,k}^{{\text{b}}}\triangleq s_{n,k+1}$ denote the left, right, top and bottom neighbors of $x_{n,k}$. The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions. Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SP rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (53), as shown in the top of this page.
@@ V-D2 Backward Message Passing
With the messages of neighbors and $v(s_{n,k})$ and SP rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
@@
Using SP rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto$ $\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}\mathrm{d}\gamma_{n,k}\mathrm{d}s_{n,k},$
@@
where $(a)$ is obtained by utilizing the property of delta function. Similarly, $m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{2}}(\alpha_{k}^{2})\propto\alpha_{k}^{2}\exp(-\alpha_{k}^{2}(\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}))$. As a result, the belief of $\alpha_{k}^{1}$ is given by
${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (57)
where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. According to SP rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})=\pi^{\mathrm{in}}_{n,j}{b}(\alpha^{1}_{k})+(1-\pi^{\mathrm{in}}_{n,j}){b}(\alpha^{2}_{k}).$ (58)
@@
$b(\gamma_{n,k})\propto m_{f_{\gamma_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}(\gamma_{n,k})$ (59)
@@
$m_{f_{x_{n,k}}\rightarrow x_{n,k}}(x_{n,k})\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (60)